OpenStudy (beccab003):

Can someone check my answer? Medal to best answer! Thanks! It will be posted below.

OpenStudy (beccab003):

Question: Simplify the following question: \[\sqrt{32x^4y^2}\] My answer: \[4x^2y \sqrt{2}\] OR \[2x^3y \sqrt{2}\]

OpenStudy (cggurumanjunath):

a

OpenStudy (cggurumanjunath):



OpenStudy (beccab003):

Thanks!! You're awesom!

OpenStudy (cggurumanjunath):

:)